OpenStudy (anonymous):

write an equation in point-slope form of the line that passes through the given point and has the given slope m? (-6,6), m =3/2?

OpenStudy (radar):

y=(3/2)x+15

OpenStudy (radar):

To check 6=(3/2)(-6)+15 6=-9+15 6=6 checks.

OpenStudy (anonymous):

so that would make it y-6=3/2(x+6) or??

OpenStudy (radar):

That would make the equation when in slope intercept form look like: y = (3/2)x+15

OpenStudy (anonymous):

oh ok thanks for helping

OpenStudy (radar):

Do you have any further questions?

OpenStudy (anonymous):

nope not for this one

OpenStudy (radar):

good luck with them.